OpenStudy (anonymous):

Pleeaassee Help!! [: Consumption of beef in United Stated increased from 9,529,000 lbs in 1950 to 27,211,000 lbs in 2000. What was the percent of increase? Round to the nearest percent. P L E A S E H E L P !

OpenStudy (anonymous):

first find the actual increase, then divide the actual increase by the original number, namely 9,529,000 then move the decimal two places right to turn your answer in to a percent

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

actually i think that is not right

OpenStudy (anonymous):

Thank You guys!

OpenStudy (anonymous):

27,211,000 - 9,529,000 = 17682000

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

((present - past)/(past)) x 100 ((27,211,000 - 9,529,000)) / 9,529,000) x 100

OpenStudy (anonymous):

\[(difference/actual \times) \times 100\]

OpenStudy (anonymous):

Answer is 185.56%

OpenStudy (anonymous):

my bad. it's late haha

OpenStudy (anonymous):

Or 186% lol

OpenStudy (anonymous):

\[1768200\div 9,529,000=1.85559...\] so your answer is \[186\%\] rounded

OpenStudy (anonymous):

THANKS GUYS!! I APPRECIATE IT!! :D

OpenStudy (anonymous):

No prob.